OpenStudy (anonymous):

(-3y2 - 7y - 9) - (4y2 + 6y + 9)

OpenStudy (anonymous):

-7y^2 -13y -18

OpenStudy (anonymous):

did that make sense ?

OpenStudy (anonymous):

Im trying to figure it out.

OpenStudy (anonymous):

you just add them up -3y^2-4y^2= -7y^2 -7y-6y=-13y -9-9=-18 -7y^2-13y-18

OpenStudy (anonymous):

I see now, but wouldn'y the 7 be positive since the 4 was a higher number?

OpenStudy (anonymous):

Or, is the four negitive?

OpenStudy (anonymous):

you have -3 and -4 if you add them together the ans will be still negative

OpenStudy (anonymous):

Okay thankyou

OpenStudy (anonymous):

yeah you can that (4y2 + 6y + 9) has a minus before it that means they all negative

OpenStudy (anonymous):

What is the measure of the angle formed by the minute hand and the hour hand of a clock when the clock shows 4:35? would this be 120 degrees